FIND THE SLOPE AND PROVE THAT THESE TWO LEGS ARE NOT PARALLEL
A. ( 3,6 )
T ( 6,2 )

Answer by Alan3354(69443)

FIND THE SLOPE AND PROVE THAT THESE TWO LEGS ARE NOT PARALLEL
A. ( 3,6 )
T ( 6,2 )
===============
What 2 legs? You only posted 2 points.
The slope of AT = diffy/diffx = (6-2)/(3-6) = -4/3
---------
PS Don't put spaces in the points --> (3,6) and (6,2)
Don't use all CAPS, either.